eleven decreased by half a number, n, is equal to seven. What does "half a number" mean and what is the number?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

eleven decreased by half a number, n, is equal to seven.
----
11 - (n/2) = 7
----
n/2 = 4
n = 8
=============
Cheers,
Stan H.

Answer by mananth(16946)

let number be x
half a number is x/2
11-x/2 = 7
multiply equation by 2
22-x=14
x=22-14
x=8